Jenny is three times as old as Sarah. In 12 years Jenny will be twice as old as Sarah is in 12 years. How old is each now?
Answer by flame8855(424)

j = 3s
j+12 = 2(s+12)
j+12 = 2s+24
3s+12 = 2s +24
s =12
j = 36